Toyota Corp.'s stock price has a variance of returns equal to 0.0295. Honda Corp.'s stock price has a variance of returns equal to 0.0505. The covariance between Toyota and Honda is 0.0675. What is the standard deviation of a portfolio consisting of 50% Toyota and 50% Honda?
Answer #1

Portfolio Standard Deviation = \sqrt{} [(WA^2*VarA) + (WB^2*VarB) + (2*WA*WB*CovAB)]

where

WA - Weight of Toyota stock =.50

WB - Weight of Honda stock =.50

VarA - Variance of Toyota stock = 0.0295

VarB - variance of Honda stock = 0.0505

CovAB - Co variance = 0.0675

Portfolio Standard Deviation = \sqrt{}[(.5^2*.0295) + (.5^2*.0505) + (2*.5*.5*.0675)]

= \sqrt{}( 0.007375 + 0.012625 + 0.03375)

= \sqrt{}0.05375

= 0.23184046238

= 23.18%